OpenStudy (anonymous):

Please help me with this problem, I can't get it.... The manager of a fish store has water that is 10% salt and water that is 25% salt. He needs to fill an aquarium with 5 gallons of water that is 20% salt. a. Write a system of linear equations that you can use to determine how many gallons of each type of salt water the manager should combine. Be sure to define your variables. b. Solve the system and determine how many gallons of each type of salt water the manager should combine. Show all your work. Please help me, thanks!

OpenStudy (jack1):

10 percent = 0.1 25 percent = 0.25 20 percent = 0.20 a) x = gallons of water that is 10% salt y = gallons of water that is 25% salt x + y = 5 gallons and 0.1(x) + 0.25(y) = 0.2 (5) now use simultaneous equations to solve for x and y

OpenStudy (jack1):

does that make sense @ka1tlyn ?

OpenStudy (anonymous):

Yes, that does make sense, thank you! Can you help my finish it too please?

OpenStudy (anonymous):

@Jack1

OpenStudy (ranga):

x + y = 5 ------ (1) 0.1x + 0.25y = 1 -------(2) Eliminate x. Multiply equation (1) by -0.1 and add it to (2) What do you get?